OpenStudy (anonymous):

x + y = 6 2x - y = 3 The system of equations a.has no solution b.has one solution c.is coincident

OpenStudy (anonymous):

Add the equations together.

OpenStudy (anonymous):

You get 3x = 9 x =3 Input your x value into the first equation. 3 + y = 6 y = 3 you will have one set of solutions.

OpenStudy (anonymous):

what about this one? 5x + 2y = 5 3x - y = 14 Which of the following equations is not a result of solving the system of equations by addition? a.11x = 19 b.11y = -55 c.11x = 33

OpenStudy (anonymous):

5x + 2y = 5 3x - y = 14 5x + 2y = 5 6x -2y = 28 11x = 33 now if you wanted to solve for y. 15x + 6y = 15 -15x + 5y = -70 11y = -55 Option A, is incorrect.

OpenStudy (anonymous):

okay and i got one last one 3x + 2y = 6 4x + y = 1 Solve the system of equations. a.(-4/5, 21/5) b.(8/11, -21/11) c.no solution

OpenStudy (anonymous):

y= -4x + 1 3x + 2(-4x+1) = 6 3x -8x + 2 = 6 -5x = 4 x = -4/5 option A

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Could you click on best response? I would appreciate it.